OpenStudy (anonymous):

The volume of water in a bowl is given by V = 1/3(pi)(h)^2(60 - h) cm^3, where h is the depth of the water in centimeters. If the depth is increasing at the rate of 4 cm/sec when the water is 10 cm deep, how fast is the volume increasing at the instant?

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

factor out V = 1/3(pi)(h)^2(60 - h) so = 20pih^2 - (1/3)pi h^3 then take derivative of that...

OpenStudy (anonymous):

dV/dt = (dV/dh)(dh/dt) need to take derivatives. If not then find the volume V at h=10 cm and then at 0.01 s later, V' for h=10.04cm, and take the numerical rate of change dV/dt= (V'-V))/0.01s

OpenStudy (anonymous):

20(pi)(h)^2-(pi)(h)^3/3

jigglypuff314 (jigglypuff314):

take the derivative of that :)

OpenStudy (anonymous):

40(pi)(h)-(pi)(h)^2

jigglypuff314 (jigglypuff314):

close :) don't forget that when taking derivatives based on time you have to include dh/dt at the ends like 40(pi)(h)(dh/dt) - (pi)(h)^2(dh/dt) before you didn't have to do that because when taking the derivative d/dx it came out as dx/dx = 1 but now since your doing it as d/dt you'll have to keep the derivative parts :)

jigglypuff314 (jigglypuff314):

you were given that dh/dt = 4 and h = 10 so plug in so that V' = 40(pi)(10)(4) - (3/2)(pi)(10)^2(4) =

OpenStudy (anonymous):

1000(pi)

jigglypuff314 (jigglypuff314):

yup :) that's what I got :)

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

1000/pie was incorrect

OpenStudy (anonymous):

i did a txt this time

jigglypuff314 (jigglypuff314):

it's not 1000/pi it's 1000pi

jigglypuff314 (jigglypuff314):

as for the yahoo answers "MechEng2030" added an extra 0 to 200liters "Michael" got what we got :) and "Keith A" didn't convert radius to cm

OpenStudy (anonymous):

we need cm^3/day

jigglypuff314 (jigglypuff314):

200 liters/day = 200000 cm^3/day

jigglypuff314 (jigglypuff314):

oh wait, do you mean for this question?

OpenStudy (anonymous):

no, actually 1000p is not it either

OpenStudy (whpalmer4):

I get \(1200\pi\) when I do the arithmetic

OpenStudy (whpalmer4):

\[\frac{dV}{dh} = 40\pi h - \pi h^2\]\[\frac{dh}{dt} = 4\]\[\frac{dV}{dt} = \frac{dV}{dh}*\frac{dh}{dt} = 4(40\pi h - \pi h^2) = 4*(40*10\pi - 10*10*pi) = 4(300\pi)\]

jigglypuff314 (jigglypuff314):

did you say that for this problem she wanted cm^3/day and Not cm^3/sec ?

OpenStudy (anonymous):

cm^3/sec

jigglypuff314 (jigglypuff314):

oh yeah, I see my mistake @whpalmer4 was right about the 1200 cm^3/day :)

OpenStudy (whpalmer4):

Answer to the original problem is \(1200\pi\).

OpenStudy (anonymous):

incorrect

OpenStudy (whpalmer4):

No, it is correct. 1200pi cm^3/s is the rate of change of the volume when the water is 10cm deep and rising 4 cm/s. If the answer your system expects is different, it is incorrect, or the problem has been transmitted incorrectly.

OpenStudy (anonymous):

my head.........

jigglypuff314 (jigglypuff314):

basically, once again we agree that 1200 cm^3/day is correct while you once again tell us that your teacher says that it is incorrect >,<

OpenStudy (anonymous):

1200 just?

OpenStudy (whpalmer4):

no, \(1200\pi \approx 3769.9111843077518862\)

jigglypuff314 (jigglypuff314):

yeah sorry, typo :P

OpenStudy (anonymous):

darn...

OpenStudy (whpalmer4):

Here's a check: at the precise moment where the height is 10 cm, the volume is \[V = \frac{1}{3}\pi(60-10)(10^2) \approx 5235.99\] At 1/1000 of a second later, the height has gone up by 1/1000*4 cm/s = 0.004 cm h = 10.004 cm \[V = \frac{1}{3}\pi(60-10.004)(10.004^2) \approx 5239.76\]\[\frac{dV}{dt} \approx \frac{\Delta V}{\Delta t} = \frac{5239.76-5235.99}{0.001} = 3770.41\] \[1200\pi \approx 3769.91\]\[\text{error} = \frac{3770.41-3769.91}{3769.91} = 0.013\%\]

jigglypuff314 (jigglypuff314):

wow nice @whpalmer4 !!!

OpenStudy (whpalmer4):

Even after all these years of doing related rates problems, I still like to do such a calculation to check my work :-)

OpenStudy (anonymous):

ughhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhh

jigglypuff314 (jigglypuff314):

mmm just stick with these answers and when possible, have your teacher give you her explanations for these problems then come back aye? :P